# Multipoint distributions of the KPZ fixed point with compactly supported initial conditions

### v2

@@ 2.1.1 The characteristic function $\chi_{\mathfrak{h}}$
The dependency on the initial condition of the entire formula is through the function $\chi_{\mathfrak{h}}$ defined on $\mathbb{C}_{\mathrm{R}}\times\mathbb{C}_{L}$. Recall that we always use $\xi$ and $\eta$ to denote a variable on the $\Gamma$-contours on $\mathbb{C}_{\mathrm{L}}$ and $\mathbb{C}_{\mathrm{R}}$ respectively throughout the paper. Note that $\chi_{\mathfrak{h}}$ is a function on $((\Gamma_{1,\mathrm{L}}\cup\Gamma_{1,\mathrm{R}})\cap\mathrm{S}_{2})\times((\Gamma_{1,\mathrm{L}}\cup\Gamma_{1,\mathrm{R}})\cap\mathrm{S}_{1})=\Gamma_{1,\mathrm{R}}\times\Gamma_{1,\mathrm{L}}$. So we will use the notation $\chi_{\mathfrak{h}}(\eta,\xi)$ in the paper. The functions $\chi_{\mathfrak{h}}$ is defined using a Brownian motion hitting expectation as follows:
@@
Then for any $\xi\in\mathbb{C}_{\mathrm{L}}$ and $\eta\in\mathbb{C}_{\mathrm{R}}$ we define
@@
$+\int_{\mathbb{R}}\mathop{}\!\mathrm{d}s\,\mathrm{e}^{-s\xi}\cdot\mathbb{E}_{\mathbf{B}(0)=s}\left[\exp\left(\bm{\tau}_{-}\eta^{2}+\mathbf{B}(\bm{\tau}_{-})\eta\right)\mathbf{1}_{\bm{\tau}_{-}>-\infty}\right]$
$-\int_{\mathbb{R}}\mathop{}\!\mathrm{d}s\,\mathbb{E}_{\mathbf{B}(0)=s}\left[\exp\left(-\bm{\tau}_{+}\xi^{2}-\mathbf{B}(\bm{\tau}_{+})\xi\right)\mathbf{1}_{\bm{\tau}_{+}<\infty}\right]\cdot\mathbb{E}_{\mathbf{B}(0)=s}\left[\exp\left(\bm{\tau}_{-}\eta^{2}+\mathbf{B}(\bm{\tau}_{-})\eta\right)\mathbf{1}_{\bm{\tau}_{-}>-\infty}\right].$
@@
The function $\chi_{\mathfrak{h}}(\eta,\xi)$ is well-defined and analytic for $\zeta\in\mathbb{C}_{\mathrm{L}}$ and $\eta\in\mathbb{C}_{\mathrm{R}}$. Moreover if $\mathfrak{h}(\alpha)=$-$\infty$ for $|\alpha|>L$, we have
@@
We verify the statement for the first term appearing on the right-hand side of (2.9), the other two terms can be treated similarly. For the well-definedness it suffices to show that the integral over the $s$-variable converges absolutely. First we split the integral into two parts depending on whether $s\leq\mathfrak{h}(0)$ or not. For $s\leq\mathfrak{h}(0)$ clearly we have $\bm{\tau}_{+}=0$ and $\mathbf{B}(\bm{\tau}_{+})=s$, hence
@@
with the integral converging absolutely since $\mathrm{Re}(\xi)<\mathrm{Re}(\eta)$ by our assumption. For the other part note that $\mathfrak{h}\in\mathrm{UC}_{c}$ is bounded above, namely, $\mathfrak{h}(\alpha)\leq\beta$ for some $\beta\in\mathbb{R}$. Define a new stopping time
@@
then clearly we have $\bm{\sigma}_{+}\leq\bm{\tau}_{+}$. Hence $\mathbb{P}_{\mathbf{B}(0)=s}(\bm{\tau}_{+}<+\infty)\leq\mathbb{P}_{\mathbf{B}(0)=s}(\bm{\sigma}_{+}<+\infty)$. Thus
@@
Now bounding $\mathrm{e}^{-\frac{(s-\beta)^{2}}{8T}}$ by $\mathrm{e}^{-\frac{(s-\beta)^{2}}{8L}}$ for $T\in[0,L]$, we have
@@
Thus we conclude that
@@
For convenience we were using the origin 0 as the starting point of the Brownian motions in the hitting formula. The next proposition shows that this is not necessary and one can start with any point on the real line $\mathbb{R}$ and get the same characteristic function $\chi_{\mathfrak{h}}$.
@@
$+\mathrm{e}^{-\omega\xi^{2}}\int_{\mathbb{R}}\mathop{}\!\mathrm{d}s\,\mathrm{e}^{-s\xi}\cdot\mathbb{E}_{\mathbf{B}(\omega)=s}\left[\exp\left(\bm{\tau}_{-}\eta^{2}+\mathbf{B}(\bm{\tau}_{-})\eta\right)\mathbf{1}_{\bm{\tau}_{-}>-\infty}\right]$
$-\int_{\mathbb{R}}\mathop{}\!\mathrm{d}s\,\mathbb{E}_{\mathbf{B}(\omega)=s}\left[\exp\left(\bm{\tau}_{-}\eta^{2}+\mathbf{B}(\bm{\tau}_{-})\eta-\bm{\tau}_{+}\xi^{2}-\mathbf{B}(\bm{\tau}_{+})\xi\right)\mathbf{1}_{|\bm{\tau}_{\pm}|<\infty}\right],$
@@
The proof of Proposition 2.4 will be given in Section 5.2. We point out that it is purely a result about the Brownian motion but we are not able to find it in the literature.
@@ 2.2 An equivalent series expansion formula
Due to the block diagonal structure of the kernel $\mathbf{K}_{1}$ and $\mathbf{K}_{\mathfrak{h}}$, the Fredholm determinant $\mathrm{D}_{\mathfrak{h}}(z_{1},\cdots,z_{m-1})$ admits a series expansion, which we will be working with more frequently in the subsequent sections. To introduce the formula, we first introduce a few notation. Given $W=(w_{1},\cdots,w_{n})\in\mathbb{C}^{n}$ and $W^{\prime}=(w_{1}^{\prime},\cdots,w_{m}^{\prime})\in\mathbb{C}^{m}$, we denote
$W\sqcup W^{\prime}:=(w_{1},\cdots,w_{n},w_{1}^{\prime},\cdots,w_{m}^{\prime})\in\mathbb{C}^{m+n}.$ (2.18)
Assume in addition that $n=m$ and $w_{i}\neq w_{j}^{\prime}$ for all $1\leq i,j\leq n$, we denote
@@
Proposition 2.5 (Series expansion for $\mathbf{D}_{\mathfrak{h}}(z_{1},\cdots,z_{m-1})$).
@@
$\mathbf{D}_{\mathfrak{h}}(z_{1},\cdots,z_{m-1})=\sum_{\begin{subarray}{c}n_{\ell}\geq 0,\\ 1\leq\ell\leq m\end{subarray}}\frac{1}{(n_{1}!\cdots n_{m}!)^{2}}\mathrm{D}^{(\mathbf{n})}_{\mathfrak{h}}(z_{1},\cdots,z_{m-1}),$ (2.20)
where $\mathbf{n}=(n_{1},\cdots,n_{m})\in(\mathbb{Z}_{\geq 0})^{m}$, and
$\begin{split}&\mathrm{D}^{(\mathbf{n})}_{\mathfrak{h}}(z_{1},\cdots,z_{m-1})=\mathrm{D}^{(\mathbf{n})}_{\mathfrak{h}}(z_{1},\cdots,z_{m-1};(\alpha_{1},\tau_{1},\beta_{1}),\cdots,(\alpha_{m},\tau_{m},\beta_{m}))\\ &=\prod_{\ell=1}^{m-1}\left(1-z_{\ell}\right)^{n_{\ell}}\left(1-z_{\ell}^{-1}\right)^{n_{\ell+1}}\left(\prod_{\ell=1}^{m}\prod_{i_{\ell}=1}^{n_{\ell}}\int_{\Sigma_{\ell,\mathrm{L}}}\mathrm{d}\mu_{\bm{z}}(\xi_{i_{\ell}}^{(\ell)})\int_{\Sigma_{\ell,\mathrm{R}}}\mathrm{d}\mu_{\bm{z}}(\eta_{i_{\ell}}^{(\ell)})\right)\prod_{\ell=1}^{m}\prod_{i_{\ell}=1}^{n_{\ell}}\frac{\mathrm{f}_{\ell}(\xi_{i_{\ell}}^{(\ell)})}{\mathrm{f}_{\ell}(\eta_{i_{\ell}}^{(\ell)})}\\ &\quad\cdot\det\begin{bmatrix}\displaystyle\chi_{\mathfrak{h}}(\eta^{(1)}_{i},\xi^{(1)}_{j})\end{bmatrix}_{1\leq i,j\leq n_{1}}\cdot\prod_{\ell=1}^{m-1}\mathrm{C}\left(\bm{\xi}^{(\ell)}\sqcup\bm{\eta}^{(\ell+1)};\bm{\eta}^{(\ell)}\sqcup\bm{\xi}^{(\ell+1)}\right)\cdot\mathrm{C}(\bm{\xi}^{(m)};\bm{\eta}^{(m)}),\end{split}$ (2.21)
with $\bm{\xi}^{(\ell)}=(\xi_{1}^{(\ell)},\cdots,\xi_{n_{\ell}}^{(\ell)})$ and $\bm{\eta}^{(\ell)}=(\eta_{1}^{(\ell)},\cdots,\eta_{n_{\ell}}^{(\ell)})$, for $1\leq\ell\leq m$. Here
@@
We remark that (2.21) looks slightly different from the version in [Liu22, (2.27)] because we use the function $\mathrm{C}$ instead of the Vandermonde type products $\Delta$. This formula also appears in [LZ25, Proposition 3.1] when $\chi_{\mathfrak{h}}(\eta,\xi)=1/(\eta-\xi)$ for the narrow wedge initial condition. The equivalence of Proposition 2.5 and Theorem 1.2 follows from [Liu22, Proposition 2.9], see also [BL19, Lemma 4.8] and [BL21, Lemma 5.6].
@@ 3.1 Characteristic function of the initial condition
A key property of the TASEP is that the distribution of any finite set of the rightmost particles, up to a fixed label $N$, is independent of the state of particles to their left. Conversely, a TASEP model with $N$ particles can be embedded into a TASEP model with infinitely many particles, where the $N$ rightmost particles correspond to the $N$-particle system, and the states of all other particles are arbitrary. This feature will be used when we characterize the initial condition of TASEP with finitely many particles.
@@
The following characterization of the initial condition is from [Liu22], see Proposition 2.13 and the subsequent discussion in that paper for further details.
@@
One could formally extend the concept of the characteristic “function” $\mathrm{ch}_{Y}(v,u)$ to an infinite system with particles labeled on $\mathbb{Z}_{+}$ and $Y=(\ldots,y_{3},y_{2},y_{1})$ by defining $\mathrm{ch}_{Y}(v,u)$ to satisfy (3.2) for all $i\in\mathbb{Z}_{+}$. We could extend it even further to a TASEP with particle labeled on $\mathbb{Z}$, while the first condition is absorbed into the second condition by allowing $i\in\mathbb{Z}$ in (3.2). The issue for these extensions is that $\mathrm{ch}_{Y}$ is not necessarily well defined as a function because of the convergence issue.
In [Liu22], the author derived a characteristic function expressed in terms of symmetric functions for any $Y$, which is well-suited for asymptotic analysis under the step or flat initial condition. As a result, the author obtained the multipoint distribution of the KPZ fixed point for both the narrow-wedge and flat initial conditions. However, the characteristic function presented in [Liu22] is not suitable for asymptotic analysis with general initial conditions. One main contribution of this paper is the following characteristic function $\mathrm{ch}_{Y}$ defined by an expectation involving random walk hitting problems, which turns out to be suitable for asymptotic analysis. The idea is heavily inspired by the seminal work [MQR21] but the formula does not follow directly from their results.
@@
$\left|\mathbb{E}_{G_{0}=z}\left[\frac{2}{(2v+2)^{G_{\tau}+1}}\left(\frac{-v}{v+1}\right)^{\tau}\mathbf{1}_{\tau<N}\right]\right|\leq\frac{2}{|2v+2|^{z+1}}\sum_{k=0}^{N-1}\sum_{w=y_{k+1}+1}^{y_{k}-1}\mathbb{P}_{G_{0}=z}[\tau=k,G_{\tau}=w]\leq\frac{1}{|2v+2|^{z}}.$ (3.7)
@@
Next we verify that the right-hand side of (3.5) satisfies (3.2) for all $1\leq i\leq N$. A taylor expansion of $u^{-i}$ at 1-1 gives
@@
Interchanging the contour integration and the expectation (which is justified by (3.7)), the above is equivalent to
$2^{z}\cdot\mathbb{E}_{G_{0}=z}\left[\oint_{0}\frac{\mathop{}\!\mathrm{d}v}{2\pi\mathrm{i}}\frac{(v+1)^{y_{i}+i-1-G_{\tau}-\tau}}{(-v)^{i-\tau}}\frac{1}{2^{G_{\tau}}}\mathbf{1}_{\tau<N}\right]=-\mathbf{1}_{z\geq y_{i}+i}{z-y_{i}-1\choose i-1}$,
which is, using the assumption that $i\leq N$,
@@
When $z\leq y_{1}$, we define $\hat{G}_{k}=G_{k+1}$ and $\hat{y}_{k}=y_{k+1}$ for $k=0,1,\cdots$. Then we have, by induction,
@@
here $\hat{\tau}:=\tau-1$. Thus we have
@@
Using the Markov property for the left-hand side of (3.10), we have

### v3

@@ 2.1.1 The characteristic function $\chi_{\mathfrak{h}}$
The dependency on the initial condition of the entire formula is through the function $\chi_{\mathfrak{h}}$ defined on $\mathbb{C}_{\mathrm{R}}\times\mathbb{C}_{L}$. Recall that we always use $\xi$ and $\eta$ to denote a variable on the $\Gamma$-contours on $\mathbb{C}_{\mathrm{L}}$ and $\mathbb{C}_{\mathrm{R}}$ respectively throughout the paper. Note that $\chi_{\mathfrak{h}}$ is a function on $((\Gamma_{1,\mathrm{L}}\cup\Gamma_{1,\mathrm{R}})\cap\mathrm{S}_{2})\times((\Gamma_{1,\mathrm{L}}\cup\Gamma_{1,\mathrm{R}})\cap\mathrm{S}_{1})=\Gamma_{1,\mathrm{R}}\times\Gamma_{1,\mathrm{L}}$. So we will use the notation $\chi_{\mathfrak{h}}(\eta,\xi)$ in the paper. The function $\chi_{\mathfrak{h}}$ is defined using a Brownian motion hitting expectation as follows:
@@
Then for any $\xi\in\mathbb{C}_{\mathrm{L}}$ and $\eta\in\mathbb{C}_{\mathrm{R}}$, we define
@@
$\quad+\int_{\mathbb{R}}\mathop{}\!\mathrm{d}s\,\mathrm{e}^{-s\xi}\cdot\mathbb{E}_{\mathbf{B}(0)=s}\left[\exp\left(\bm{\tau}_{-}\eta^{2}+\mathbf{B}(\bm{\tau}_{-})\eta\right)\mathbf{1}_{\bm{\tau}_{-}>-\infty}\right]$
$\quad-\int_{\mathbb{R}}\mathop{}\!\mathrm{d}s\,\mathbb{E}_{\mathbf{B}(0)=s}\left[\exp\left(-\bm{\tau}_{+}\xi^{2}-\mathbf{B}(\bm{\tau}_{+})\xi\right)\mathbf{1}_{\bm{\tau}_{+}<\infty}\right]\cdot\mathbb{E}_{\mathbf{B}(0)=s}\left[\exp\left(\bm{\tau}_{-}\eta^{2}+\mathbf{B}(\bm{\tau}_{-})\eta\right)\mathbf{1}_{\bm{\tau}_{-}>-\infty}\right].$
@@
For $\mathfrak{h}\in\mathrm{UC}_{c}$, the function $\chi_{\mathfrak{h}}(\eta,\xi)$ is well-defined and analytic for $\zeta\in\mathbb{C}_{\mathrm{L}}$ and $\eta\in\mathbb{C}_{\mathrm{R}}$. Moreover, if $\mathfrak{h}(\alpha)=$-$\infty$ for $|\alpha|>L$, we have
@@
We verify the statement for the first term appearing on the right-hand side of (2.9), the other two terms can be treated similarly. For the well-definedness it suffices to show that the integral over the $s$-variable converges absolutely. Take $L>0$ large enough such that $\mathrm{supp}(\mathfrak{h})\subset[-L,L]$. We split the integral into two parts depending on whether $s\leq\mathfrak{h}(0)$ or not. For $s\leq\mathfrak{h}(0)$ clearly we have $\bm{\tau}_{+}=0$ and $\mathbf{B}(\bm{\tau}_{+})=s$, hence
@@
with the integral converging absolutely since $\mathrm{Re}(\xi)<\mathrm{Re}(\eta)$ by our assumption. For the other part, note that $\mathfrak{h}\in\mathrm{UC}_{c}$ is bounded above, namely, $\mathfrak{h}(\alpha)\leq\beta$ for some $\beta\in\mathbb{R}$. Define a new stopping time
@@
then clearly we have $\bm{\sigma}_{+}\leq\bm{\tau}_{+}$. Hence $\mathbb{P}_{\mathbf{B}(0)=s}(\bm{\tau}_{+}<+\infty)\leq\mathbb{P}_{\mathbf{B}(0)=s}(\bm{\sigma}_{+}<+\infty)$. Thus,
@@
Now by bounding $\mathrm{e}^{-\frac{(s-\beta)^{2}}{8T}}$ by $\mathrm{e}^{-\frac{(s-\beta)^{2}}{8L}}$ for $T\in[0,L]$, we have
@@
Thus, we conclude that
@@
For convenience, we were using the origin 0 as the starting point of the Brownian motions in the hitting formula. The next proposition shows that this is not necessary, and one can start with any point on the real line $\mathbb{R}$ and get the same characteristic function $\chi_{\mathfrak{h}}$.
@@
$\quad+\mathrm{e}^{-\omega\xi^{2}}\int_{\mathbb{R}}\mathop{}\!\mathrm{d}s\,\mathrm{e}^{-s\xi}\cdot\mathbb{E}_{\mathbf{B}(\omega)=s}\left[\exp\left(\bm{\tau}_{-}\eta^{2}+\mathbf{B}(\bm{\tau}_{-})\eta\right)\mathbf{1}_{\bm{\tau}_{-}>-\infty}\right]$
$\quad-\int_{\mathbb{R}}\mathop{}\!\mathrm{d}s\,\mathbb{E}_{\mathbf{B}(\omega)=s}\left[\exp\left(\bm{\tau}_{-}\eta^{2}+\mathbf{B}(\bm{\tau}_{-})\eta-\bm{\tau}_{+}\xi^{2}-\mathbf{B}(\bm{\tau}_{+})\xi\right)\mathbf{1}_{|\bm{\tau}_{\pm}|<\infty}\right],$
@@
The proof of Proposition 2.4 will be given in Section 5.2. We point out that it is purely a result about the Brownian motion, but we are not able to find it in the literature.
@@ 2.2 An equivalent series expansion formula
Due to the block diagonal structure of the kernel $\mathbf{K}_{1}$ and $\mathbf{K}_{\mathfrak{h}}$, the Fredholm determinant $\mathrm{D}_{\mathfrak{h}}(z_{1},\ldots,z_{m-1})$ admits a series expansion, which we will be working with more frequently in the subsequent sections. To introduce the formula, we first introduce a few notations. Given $W=(w_{1},\ldots,w_{n})\in\mathbb{C}^{n}$ and $W^{\prime}=(w_{1}^{\prime},\ldots,w_{m}^{\prime})\in\mathbb{C}^{m}$, we denote
$W\sqcup W^{\prime}:=(w_{1},\ldots,w_{n},w_{1}^{\prime},\ldots,w_{m}^{\prime})\in\mathbb{C}^{m+n}.$ (2.18)
Assume in addition that $n=m$ and $w_{i}\neq w_{j}^{\prime}$ for all $1\leq i,j\leq n$. We denote
@@
Proposition 2.5 (Series expansion for $\mathbf{D}_{\mathfrak{h}}(z_{1},\ldots,z_{m-1})$).
@@
$\mathbf{D}_{\mathfrak{h}}(z_{1},\ldots,z_{m-1})=\sum_{\begin{subarray}{c}n_{\ell}\geq 0,\\ 1\leq\ell\leq m\end{subarray}}\frac{1}{(n_{1}!\cdots n_{m}!)^{2}}\mathrm{D}^{(\mathbf{n})}_{\mathfrak{h}}(z_{1},\ldots,z_{m-1}),$ (2.20)
where $\mathbf{n}=(n_{1},\ldots,n_{m})\in(\mathbb{Z}_{\geq 0})^{m}$, and
$\begin{split}&\mathrm{D}^{(\mathbf{n})}_{\mathfrak{h}}(z_{1},\ldots,z_{m-1})=\mathrm{D}^{(\mathbf{n})}_{\mathfrak{h}}(z_{1},\ldots,z_{m-1};(\alpha_{1},\tau_{1},\beta_{1}),\ldots,(\alpha_{m},\tau_{m},\beta_{m}))\\ &=\prod_{\ell=1}^{m-1}\left(1-z_{\ell}\right)^{n_{\ell}}\left(1-z_{\ell}^{-1}\right)^{n_{\ell+1}}\left(\prod_{\ell=1}^{m}\prod_{i_{\ell}=1}^{n_{\ell}}\int_{\Gamma_{\ell,\mathrm{L}}}\mathrm{d}\mu_{\bm{z}}(\xi_{i_{\ell}}^{(\ell)})\int_{\Gamma_{\ell,\mathrm{R}}}\mathrm{d}\mu_{\bm{z}}(\eta_{i_{\ell}}^{(\ell)})\right)\prod_{\ell=1}^{m}\prod_{i_{\ell}=1}^{n_{\ell}}\frac{\mathrm{f}_{\ell}(\xi_{i_{\ell}}^{(\ell)})}{\mathrm{f}_{\ell}(\eta_{i_{\ell}}^{(\ell)})}\\ &\quad\cdot\det\begin{bmatrix}\displaystyle\chi_{\mathfrak{h}}(\eta^{(1)}_{i},\xi^{(1)}_{j})\end{bmatrix}_{1\leq i,j\leq n_{1}}\cdot\prod_{\ell=1}^{m-1}\mathrm{C}\left(\bm{\xi}^{(\ell)}\sqcup\bm{\eta}^{(\ell+1)};\bm{\eta}^{(\ell)}\sqcup\bm{\xi}^{(\ell+1)}\right)\cdot\mathrm{C}(\bm{\xi}^{(m)};\bm{\eta}^{(m)}),\end{split}$ (2.21)
with $\bm{\xi}^{(\ell)}=(\xi_{1}^{(\ell)},\ldots,\xi_{n_{\ell}}^{(\ell)})$ and $\bm{\eta}^{(\ell)}=(\eta_{1}^{(\ell)},\ldots,\eta_{n_{\ell}}^{(\ell)})$, for $1\leq\ell\leq m$. Here
@@
We remark that (2.21) looks slightly different from the version in [Liu22, (2.27)] because we use the function $\mathrm{C}$ instead of the Vandermonde type products $\Delta$. The formula (2.21) also appears in [LZ25, Proposition 3.1] when $\chi_{\mathfrak{h}}(\eta,\xi)=1/(\eta-\xi)$ for the narrow wedge initial condition. The equivalence of Proposition 2.5 and Theorem 1.2 follows from [Liu22, Proposition 2.9], see also [BL19, Lemma 4.8] and [BL21, Lemma 5.6].
@@ 3.1 Characteristic function of the initial condition
One feature of the TASEP is that the distribution of any finite set of the rightmost particles, up to a fixed label $N$, is independent of the state of particles to their left. Conversely, a TASEP model with $N$ particles can be embedded into a TASEP model with infinitely many particles, where the $N$ rightmost particles correspond to the $N$-particle system, and the states of all other particles are arbitrary. This feature will be used when we characterize the initial condition of TASEP with finitely many particles.
@@
The following characterization of the initial condition comes from [Liu22], see Proposition 2.13 and the subsequent discussion in that paper for further details.
@@
One could formally extend the concept of the characteristic “function” $\mathrm{ch}_{Y}(v,u)$ to an infinite system with particles labeled on $\mathbb{Z}_{+}$ and $Y=(\ldots,y_{3},y_{2},y_{1})$ by defining $\mathrm{ch}_{Y}(v,u)$ to satisfy (3.2) for all $i\in\mathbb{Z}_{+}$. We could extend it even further to a TASEP with particle labeled on $\mathbb{Z}$, while the first condition is absorbed into the second condition by allowing $i\in\mathbb{Z}$ in (3.2). The issue for these extensions is that $\mathrm{ch}_{Y}$ is not necessarily well-defined as a function because of the convergence issue.
In [Liu22], the author derived a characteristic function expressed in terms of symmetric functions for any $Y$, which is well-suited for asymptotic analysis under the step or flat initial condition. As a result, the author obtained the multipoint distribution of the KPZ fixed point for both the narrow-wedge and flat initial conditions. However, the characteristic function presented in [Liu22] is not suitable for asymptotic analysis with general initial conditions. One main contribution of this paper is the following characteristic function $\mathrm{ch}_{Y}$ defined by an expectation involving random walk hitting problems, which turns out to be suitable for asymptotic analysis. The idea is heavily inspired by the seminal work [MQR21]: we compared the two formulas of [MQR21] and [Liu22] and guessed the identity. We remark that a similar hitting expectation kernel expression has also been derived in [BLSZ23], which generalized the approach of [MQR21]. Another interesting connection is that the left-hand side of the equation (3.9) in our proof has a similar structure with an expression of the $G(z_{1},z_{2})$ function in [BLSZ23, Proposition 4.6], although they are still different [^3] The left-hand side of (3.9) has two factors $2^{z}$ and $2^{-G_{\tau}}$ which the $G(z_{1},z_{2})$ function in [BLSZ23] does not have. This results in the different meanings of these two quantities: (3.9) is a binomial coefficient, while $G(z_{1},z_{2})$ in [BLSZ23] is a probability. .
@@
$\begin{split}\left|\mathbb{E}_{G_{0}=z}\left[\frac{2}{(2v+2)^{G_{\tau}+1}}\left(\frac{-v}{v+1}\right)^{\tau}\mathbf{1}_{\tau<N}\right]\right|&\leq\frac{2}{|2v+2|^{z+1}}\sum_{k=0}^{N-1}\sum_{w=y_{k+1}+1}^{y_{k}-1}\mathbb{P}_{G_{0}=z}[\tau=k,G_{\tau}=w]\\ &\leq\frac{1}{|2v+2|^{z}}.\end{split}$ (3.7)
@@
Next, we verify that the right-hand side of (3.5) satisfies (3.2) for all $1\leq i\leq N$. A Taylor expansion of $u^{-i}$ at 1-1 gives
@@
By interchanging the contour integration and the expectation (which is justified by (3.7)), the above is equivalent to
$2^{z}\cdot\mathbb{E}_{G_{0}=z}\left[\oint_{0}\frac{\mathop{}\!\mathrm{d}v}{2\pi\mathrm{i}}\frac{(v+1)^{y_{i}+i-1-G_{\tau}-\tau}}{(-v)^{i-\tau}}\frac{1}{2^{G_{\tau}}}\mathbf{1}_{\tau<N}\right]=-\mathbf{1}_{z\geq y_{i}+i}{z-y_{i}-1\choose i-1}$, (3.9)
which is, by using the assumption that $i\leq N$ and the fact that the $v$-integral vanishes when $i\geq\tau$,
@@
When $z\leq y_{1}$, we define $\hat{G}_{k}=G_{k+1}$ and $\hat{y}_{k}=y_{k+1}$ for $k=0,1,\ldots$. Then we have, by induction,
@@
here $\hat{\tau}:=\tau-1$. Thus, we have
@@
By using the Markov property for the left-hand side of (3.11), we have